Verify that the given differential equation is an exact: 

(y2-1)dx+(2xy+3y)dy=0


1
Expert's answer
2022-01-06T12:05:30-0500
"(y^2-1)dx+(2xy+3y)dy=0"

"P(x, y)=y^2-1, \dfrac{\partial P}{\partial y}=2y"

"Q(x, y)=2xy+3y, \dfrac{\partial Q}{\partial x}=2y"

 Use the test for exactness:


"\dfrac{\partial Q}{\partial x}= \dfrac{\partial P}{\partial y}"

Since


"\dfrac{\partial Q}{\partial x}=2y= \dfrac{\partial P}{\partial y}"

the differential equation

"(y^2-1)dx+(2xy+3y)dy=0"

is an exact equation.